Write the quadratic equation in general form
(x-3)^2=0

Answer by atique.ah(27)

Solution:
simply use the formula (a+b)^2=a^2+b^2+2ab
or simply (x-3)x (x-3)=x^2-6x+9=0
where a=1,b=-6,c=9